OpenStudy (anonymous):

Solve for x. 6 – 3|2x + 6| = 0

OpenStudy (justnick09):

x=4

OpenStudy (anonymous):

2x² - x - 6 = 0 2 * -6 = -12 -4 * 3 = -12 -4 + 3 = -1 2x² - 4x + 3x - 6 = 0 (same as original) (2x² - 4x) + (3x - 6) = 0 2x(x - 2) + 3(x - 2) = 0 (2x + 3)(x - 2) = 0 2x + 3 = 0 2x = -3 x = -3/2 x - 2 = 0 x = 2 x = -3/2 and x = 2

OpenStudy (justnick09):

that's not correct at all^